DX LUCI AO/Commissioning

06:16
  • Waiting for seeing limited observations to complete.
6:33
  • Preseting to a 16 mag quasar.
  • 15.49 catalog brightness.
  • Bin 4, 200 Hz
  • Flao magnitude shows 15.8
  • Autogains too high (At maximum of all ranges, or nearly so).
  • Gains I selected: 2.00, 1.35, 1.00
06:46
  • Bizarre error on the instrument manager side, the preset repeated
  • Datasets: 065022, 065204, 065224, 065243
07:06
  • Jochen complains about elongated stars. I set TT up to 3.50, what the Autogain found.
  • There is quite a bit of vibration in OVMS.
  • Dataset: 070702
  • Dataset: 072510
  • Worried about forces along damaged area, setting TT gain to 3.00
  • Dataset: 072632
  • Just noticed, we're actually at 100 Hz.
08:07
  • Xianyu's fix really seems to have helped. We haven't had any offset errors over the past two nights.
  • There have been a (very, very) few SafeSkip frames. (122 over a couple of hours)
08:44
  • A few more Safe Skp frames occurred.
  • The seeing is extremely good, so I've lowered the TT gain back to 2.00
  • I suspect that it was over-driving. I lowered HO1 to 1.00 and HO2 to 0.90
  • High current warning from #466 and #338.
  • These are among the "Usual Suspects".
08:53
  • Loop opened due to high forces.
  • It was as though the forces were continually building up on those actuators until the loop opened. After re-closing the loop, they haven't given any errors.
  • Just before the loop opened, there was a bit of a checker-board look on the current plot in the mirror GUI. Almost as though the mirror was trying to drive a waffle mode.
  • The seeing has been spectacularly good. 0.8 is the worst that I've seen.
09:13
  • #376 is showing high currents (0.8)
  • Position measurements for this actuator isn't different from its neighbors.
09:46
  • I'm going to remove 466 and 338 from the ICC list (added to no icc)
  • safe skip made the loop open again.
10:13
  • It looks like removing 338 and 466 from ICC didn't help.
  • Preset to a blazar.
  • Bin 3, 200Hz
  • Gain 1.94 0.99 0.60
  • 14.7 mag
  • Dataset: 101534, 101609, 101633, 101644
  • I've reverted the no ICC list, changing it didn't seem to help.
  • next time the system is restarted, it will revert.
11:00
  • GCS moved the star off of the acquisition frame.
  • Need a new preset.
  • Going to a PSF star
  • Hopefully, it'll be bin 3
  • Skipping "Check flux" to force it to stay in bin3, it was borderline
11:22
  • Jochen didn't like the previous one, it was too faint.
  • Trying a new one.
  • FLAO is showing 15.4, catalog 14.9
  • Bin 4, 100 Hz
11:40
  • Preset back to the Blazar.
12:17
  • New preset, PSF star
  • A very red one, to make it bright in K.
  • Actuator 376 is showing very high forces.
  • Looks to be staying in bin 3.
  • Catalog: mag 14.46
  • Estimated 14.8
  • Same gains as before, to try to get similar AO performance, to make it a good PSF star.
  • Dataset: 122935, 123003, 123039, 123107 (during offset), 123135
12:53
  • There was the signature of heavy vibrations in Jochen's images, causing double images. I increased the TT gain.
  • I did notice, however, that there was a lot of vibration in the OVMS plot. (around 12:45)
  • Now we are complaining about a lot of trefoil.
13:00
  • We went to one of the PSF stars, by mistake.
  • Looks to be going to bin 4, this time.
  • I'm seeing mag 15.2
13:15
  • Continuing to see if we can find the source of the double images.
  • The OVMS vibrations are not unusually high. The TT gain is not over driving,
  • It could be something to do with passive flexure compensation.
13:16
  • The processes on both W-units show that they have disconnected from msgdb.
  • The AO loop is working fine, however.
13:21
  • Now the SX AdSec processes have disconnected.
13:28
  • Background is getting higher, now my total mag is 14.8
13:36
  • Stopping due to twilight.

-- %USERSIG{GregoryTaylor - 2017-12-30}%
Topic revision: r2 - 30 Dec 2017, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback